Problem 18-07
Refunding Analysis

Mullet Technologies is considering whether or not to refund a $175 million, 15% coupon, 30-year bond issue that was sold 5 years ago. It is amortizing $3 million of flotation costs on the 15% bonds over the issue's 30-year life. Mullet's investment banks have indicated that the company could sell a new 25-year issue at an interest rate of 10% in today's market. Neither they nor Mullet's management anticipate that interest rates will fall below 10% any time soon, but there is a chance that rates will increase.

A call premium of 14% would be required to retire the old bonds, and flotation costs on the new issue would amount to $7 million. Mullet's marginal federal-plus-state tax rate is 40%. The new bonds would be issued 1 month before the old bonds are called, with the proceeds being invested in short-term government securities returning 6% annually during the interim period.

  1. Conduct a complete bond refunding analysis. What is the bond refunding's NPV? Do not round intermediate calculations. Round your answer to the nearest cent.

  2. What factors would influence Mullet's decision to refund now rather than later?
Answer #1

The question was very long. I was able to complete first part of the question (calculation of bond refunding's NPV) within the available time.

_______

Step 1: Calculate Initial Investment Outlay

The value of initial investment outlay is determined as below:

Call Premium on Old Issue = Value of Old Issue*Call Premium Rate*(1-Tax Rate) = 175,000,000*14%*(1-40%) = $14,700,000

New Flotation Cost = $7,000,000

Tax Savings on Old Flotation Cost = Old Flotation Costs*Remaining Maturity/Original Maturity*Tax Rate = 3,000,000*(30 - 5)/30*40% = $1,000,000

Additional Interest on Old Issue for 1 Month = Value of Old Issue*1/12*Coupon Rate*(1-Tax Rate) = 175,000,000*1/12*15%*(1-40%) = $1,312,500

Interest Earned on Short Term Investment for 1 Month = Value of Proceeds*1/12*Return on Short Term Government Securities*(1-Tax Rate) = 175,000,000*1/12*6%*(1-40%) = $525,000

Now, we can determine initial investment outlay as follows:

Initial Investment Outlay = Call Premium on Old Issue + New Flotation Cost - Tax Savings on Old Flotation Cost + Additional Interest on Old Issue for 1 Month - Interest Earned on Short Term Investment for 1 Month = 14,700,000 + 7,000,000 - 1,000,000 + 1,312,500 - 525,000 = $21,487,500

______

Step 2: Determine Annual Cash Flows

To determine annual cash flows, we need to arrive at the value of annual flotation cost tax effects and annual interest savings due to refunding as below:

Total Amortization Tax Effects = Annual Tax Savings on New Flotation - Tax Savings Lost on Old Flotation = 7,000,000/25*40% - 3,000,000/30*40% = $72,000

Net After Tax Interest Savings = Annual After Tax Interest on Old Bond - Annual After Tax Interest on New Bond = Value of Old Issue*Coupon Rate*(1-Tax Rate) - Value of New Issue*Coupon Rate*(1-Tax Rate) = 175,000,000*15%*(1-40%) - 175,000,000*10%*(1-40%) = $5,250,000

Annual Cash Flows = Total Amortization Tax Effects + Net After Tax Interest Savings = 72,000 + 5,250,000 = $5,322,000

______

Step 2: Determine Bond Refunding's NPV

The value of bond refunding's NPV is calculated as below:

NPV of Bond Refunding's Decision = Present Value of Annual Cash Flows - Initial Investment Outlay = 5,322,000*PVA(6%,25 Years) - 21,487,500 = 5,322,000*12.783356 = 68,033,021.63 - 21,487,500 = $46,545,521.63 or $46,545,521

______

Notes

1) PVA indicates Present Value of Annuity for $1. It can be obtained with the use of present value tables.